Using prices in 2020 as a base year, calculate real GDP in 2020 and 2021.

                Q 2020    P 2020  Q 2021    P 2021

Apples      120          1        240              3

oranges    65            2        90               4

plams        80            4      120             5

(Q) represent Quantity and (P) price


1
Expert's answer
2021-02-01T09:01:54-0500

Let's calculate the real GDP in 2020 year:


"Real\ GDP_{2020}=120\cdot\$1+65\cdot\$2+80\cdot\$4=\$570."

Let's calculate the real GDP in 2021 year:


"Real\ GDP_{2021}=240\cdot\$1+90\cdot\$2+120\cdot\$4=\$900."